OpenStudy (jb1625):

Will Medal What is the value of 3 • (–8) – 2 • (–5)?

OpenStudy (defiance):

First 3 x -8 = -24

OpenStudy (defiance):

Now, 2 x -5 = -10

OpenStudy (jb1625):

Ok Thank you

OpenStudy (defiance):

Now we have (-24) - (-10)= -14

OpenStudy (defiance):

Two -'s make a +. If they're next to eachother. <Just to stop confusion>

OpenStudy (jb1625):

Thank you

OpenStudy (defiance):

No problem.